Is 1/2y=1 a linear equation?
Answer by jim_thompson5910(35256)

1/2y=1


1y=1*2


1y = 2


y = 2/1


y = 2


Since the equation is in the form y = mx+b, where m = 0 and b = 2, this means that y = 2 is a linear equation.

So by extension, is also a linear equation.